OpenStudy (anonymous):

if f(x)= sqrt (8-5x), find a formula for f^(-1)(x)

OpenStudy (anonymous):

\[x=\sqrt{8-5y}\] \[x^2=8-5y\] \[5y=8-x^2\] \[y=\frac{8-x^2}{5}\] \[f^{-1}(x)=\frac{8-x^2}{5}\]

OpenStudy (anonymous):

inverse of (3x-10)/(9x+9)